Statistics from a​ college's climate center indicate that the city the college is in gets an average of 35.9​" of rain each​ year, with a standard deviation of 4.1​". Assume that a Normal model applies. Complete parts a through d below. ​

A. During what percentage of years does this city get more than 41​" of​ rain?

The percentage of years with more than 41​" of rain is ___ ​%.
Answer #1

X ~ N ( µ = 35.9 , σ = 4.1 )
We covert this to standard normal as
P ( X < x) = P ( (Z < X - µ ) / σ )
P ( X > 41 ) = P(Z > (41 - 35.9 ) / 4.1 )
= P ( Z > 1.24 )
= 1 - P ( Z < 1.24 )
= 1 - 0.8925
= 0.1075
= 10.75 %